Status of ATF/ATF2 and machine schedule in this fall Impact of the earthquake Recovery status, alignment Schedule in this fall N. Terunuma, KEK.
Published byModified over 4 years ago
Presentation on theme: "Status of ATF/ATF2 and machine schedule in this fall Impact of the earthquake Recovery status, alignment Schedule in this fall N. Terunuma, KEK."— Presentation transcript:
Status of ATF/ATF2 and machine schedule in this fall Impact of the earthquake Recovery status, alignment Schedule in this fall N. Terunuma, KEK
Power line recovery Rough Alignment (LINAC) RF Blackout ½ power Vacuum ATF2 Shield Cooling Water (BT) (DR) Summer Shutdown Test Beam KEK Requested Standby at Home Power Distri- bution ATF Shield Radiation Safety
Impact of the earthquake on ATF Photo: 2011/March/11 15:55
Checkout of Electricity and Water in May DR magnet PW station, 6.6kV Cooling Water pipes
A test beam passes all beamline 2011/5/25 Injector 2011/5/26 LINAC 5/31 DR 1 st turn Single bunch, 0.78 Hz, 0.3 x 10 10 e/bunch DR&ATF2 5/27 BT end 6/2 DR storage 6/2 MS1X 6/3 MS1FF 6/3 MS2FF 6/3 MSPIP 2011/06/03 N.Terunuma
Stored beam in DR (x10 10 e/bunch) A stored beam was delivered to the dump of ATF2. No critical damage on the accelerator was found. 6/1 6/3 6/8 6/106/13 6/17 6/24 6/30 0.5x10 10 1.0x10 10 3 trains XSR ready Compton 2-mirror FONT 3 trains Cavity BPMs FFTB mover HAPS Freq. Adjust V-emittance (XSR) ~30pm DR rough alignment for checkout was continued in daytime. Inspection of the radiation safety
Welcome back from oversea 17 persons May - August
Status of Alignment All in the ATF building was shaken (10 cm or more?) by Great East Japan Earthquake. Linac/BT Lift up the downstream section close to the BT/DR level. It will reduce the workload of DR tuning. DR ¾ of DR has been re-aligned. East arc will be done in this week. Expanded North-South about 1.5 mm Rematch the optics ATF2: Just surveyed. Waiting the alignment work in October.
Alignment: LINAC/BT level comments Measured after earthquake Re-aligned LINAC Table Re-aligned Cavity LINAC’s drop to end and jump to BT had been known since the DR startup. It might be caused by the DR construction after LINAC alignment. It brought a troublesome tuning to us. We corrected these by using a power-recovering long shutdown after the earthquake. BT was re-aligned but plots were not yet prepared.
DR displacement (Enlarged image) after earthquake Rough alignment in May 1.5 mm wider(N-S) than the original design. When?? New design was defined as wider but same circumference.
Alignment: DR level aligned in June, aligned in summer Vertical (mm) -0.5 0.0 0.5 NorthWest ArcSouth East Arc is under realignment. Target: |x|< 90 um
Alignment: DR Horizontal (mm) -0.5 0.0 1.0 0.5 Longitudinal (mm) -0.5 0.0 1.0 0.5 aligned in summer Othersaligned in May NorthWest Arc South East Arc is under realignment. Target: |x|< 90 um
Waiting for alignment: ATF2 Floor of ATF2 sank about 1.5 mm. Level of the EXT-ATF2 magnets EXT FF Vertical (mm) QD0 mover is OFF QF1 mover is ON FFTB movers are “zero” position. Target: +-0.1 mm 0.0 -2.0 -3.0
Waiting for alignment: ATF2 Horizontal EXT FF Target: +-0.1 mm 0.0 -2.0 -3.0 0.0 -2.0 -3.0 -4.0 1.0 (mm) Longitudinal EXT FF Datum on floor We are checking the definition of the datum on the floor used at the ATF2 construction. Offset?
Waiting for alignment: ATF2 Roll of magnets (mrad) 1.0 - 1.0 0 EXT Target: +-0.1 mrad Pitch (mrad) 1.0 - 1.0 FF
DR Checkout of beam instruments EXT-FF Low emittance, Compton and Others ATF R&D ATF2 R&D Alignment DR survey ATF2 &TB meeting Recovery of 10 pm (DR) Recovery of 300 nm Goal-1 and others extraction survey
New IP-BPM installation at IP Low emittance, Compton and Others Goal-1? and others Goal-2? and others Summer Shutdown GDE meeting? JPS meeting
Alignment: DR After earthquake, aligned in June, aligned in summer - 2.0 mm +2.0 mm HorizontalLongitudinal all B Q Sx - 2.0 mm +2.0 mm
Hardware updates 1/? Pol e+ Compton: Upgrade of the 2-mirror system to 4-mirror system is underway. (Hiroshima, KEK) Old 2-mirror system in DR was replaced with a dummy chamber in August. New 4-mirror system is under testing beside the beamline. Installation will be done in the maintenance week after the full checkout of the optical system. It will be in October or later.